OpenStudy (anonymous):

-4 ( 2x + 2) = -4(x + 1)/4 Edit: I got to -8 - 8 * 4 = -4x - 4 and I'm not sure about the next step

OpenStudy (anonymous):

-4(2x+2)=-(x+1) the 4 on the right side cancels out.

OpenStudy (anonymous):

now simplify: -8x-8=-x-1

OpenStudy (anonymous):

-8x+1x=-1+8 and the rest is for you, let me know if you are stuck on this step

OpenStudy (anonymous):

I'm sorry I don't know how you got from -8x-8=-x-1 to -8x+1x=-1+8 It's been a while since I've done Math, Could you please elaborate?

OpenStudy (anonymous):

sure, these are algebraic rules, which you must know or learn all i did was move the (-x) to the left side which becomes + (note -x=-1x) and then I moved the (-8) to the right which becomes +

OpenStudy (anonymous):

ok, hope thats good for you, I am heading off now, see ya.

OpenStudy (anonymous):

Thank you for your help, I solved the problem. Seriously thank you.